The defect rate for your product has historically been about 1.50​%. For a sample size of 300​, the upper and lower 3​-sigma control chart limits​ are:

UCLP=0.0361

LCLP=

I NEED THE LOWER CONTROL LIMIT PLEASE :)
Answer #1

LCL = p - z [ sprt(p *(1 - p) / n]

Where p = 0.015

Z = 3 (3 sigma)

N = 300

LCL = 0.015 - 3 * sqrt(0.015 * (1 - 0.015) / 300)

= 0.015 - 0.021053

= - 0.006053

Do give a positive rating if you think this answer helped